# Taming OOD Actions for Offline Reinforcement Learning: An Advantage-Based Approach

### v1

@@ Abstract
Offline reinforcement learning (RL) aims to learn decision-making policies from fixed datasets without online interactions, providing a practical solution where online data collection is expensive or risky. However, offline RL often suffers from distribution shift, resulting in inaccurate evaluation and substantial overestimation on out-of-distribution (OOD) actions. To address this, existing approaches incorporate conservatism by indiscriminately discouraging all OOD actions, thereby hindering the agent’s ability to generalize and exploit beneficial ones. In this paper, we propose Advantage-based Diffusion Actor-Critic (ADAC), a novel method that systematically evaluates OOD actions using the batch-optimal value function. Based on this evaluation, ADAC defines an advantage function to modulate the Q-function update, enabling more precise assessment of OOD action quality. We design a custom PointMaze environment and collect datasets to visually reveal that advantage modulation can effectively identify and select superior OOD actions. Extensive experiments show that ADAC achieves state-of-the-art performance on almost all tasks in the D4RL benchmark, with particularly clear margins on the more challenging tasks. Our code is available at https://anonymous.4open.science/r/adac-14D0.
@@ 1 Introduction
Offline reinforcement learning (RL) [22, 23] focuses on learning decision-making policies solely from previously collected datasets, entirely circumventing the need for online interactions with the environment. This paradigm is particularly appealing for applications where online data collection is prohibitively expensive or poses safety concerns, thereby fueling its growing adoption [16, 28]. While conceptually attractive, offline RL often suffers from distribution shift between the behavior policy and the learned policy. The policy evaluation on out-of-distribution (OOD) actions is prone to extrapolation error [11], which can be amplified through bootstrapping, leading to significant overestimation [21].
To mitigate overestimation, a common strategy in offline RL is to incorporate conservatism into algorithm design. Value-based methods [21, 19, 25] achieve this by learning a pessimistic value function that reduces the estimated value of OOD actions to discourage their selection. Alternatively, policy-based methods [10, 11, 36] enforce conservatism by constraining the learned policy to remain close to the behavior policy, thereby avoiding querying OOD actions. Similarly, conditional sequence modeling approaches [7, 14, 2] inherently induce conservative behavior by restricting the policy to imitate the behavior contained in the offline dataset. In a distinct manner, model-based approaches [18, 40, 34] ensure conservatism by learning a pessimistic dynamics model where uncertainty-based penalization systematically underestimates the value of OOD actions.
@@
To this end, we propose Advantage-based Diffusion Actor-Critic (ADAC), a novel method that systematically assesses the quality of OOD actions, selectively encourages beneficial ones, and discourages risky ones to maintain conservatism. Our key idea builds on the insight that the optimal value function provides a natural baseline for evaluating state desirability. Specifically, we regard an OOD action as advantageous if it can move the current state to a successor state whose value, under the optimal value function, exceeds that of any reachable state under the behavior policy. Since the true optimal value function is inaccessible from offline data, we adopt the batch-optimal value function ($V_{\mu}^{\ast}$ in Figure 1) as a principled surrogate. We show theoretically that the batch-optimal value function can be reliably approximated via expectile regression on the dataset. Based on this approximation, we define an advantage function to assess the desirability of actions, which is then used to modulate the temporal difference (TD) target during Q-function (critic) learning. The policy (actor) is parameterized as a diffusion model due to its strong expressiveness [13, 36] and is updated under the guidance of the learned Q-function. We summarize the full architecture of ADAC in Figure 1. To the best of our knowledge, ADAC represents a pioneering attempt to systematically evaluate OOD actions, enabling the selection of beneficial ones. As a result, ADAC achieves state-of-the-art (SOTA) performance on the majority of the D4RL [9] benchmark tasks.
@@ 2 Preliminaries
Offline Reinforcement Learning. RL problems are commonly formulated within the framework of a Markov Decision Process (MDP), defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,\rho_{0},P,\gamma)$. Here, $\mathcal{S}$ denotes the state space, $\mathcal{A}$ represents the action space, $r(\bm{s},\bm{a}):\mathcal{S}\times\mathcal{A}\to[-R_{\max},R_{\max}]$ is a bounded reward function, $\rho_{0}(\bm{s})$ specifies the initial state distribution, $P(\bm{s^{\prime}}\,|\,\bm{s},\bm{a}):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}_{+}$ defines the transition dynamics, and $\gamma\in(0,1)$ is the discount factor [35].
@@
Diffusion Model. Diffusion models [32, 13, 33] consist of a forward process that corrupts data with noise and a reverse process that reconstructs data from noise. Specifically, the forward process is conducted by gradually adding Gaussian noise to samples $\bm{x_{0}}$ from an unknown data distribution $p_{\theta}(\bm{x_{0}})$, formulated as:
@@
where $T$ denotes the total number of diffusion steps, and $\alpha_{t}$ controls the variance of the added noise at each step $t$. The reverse process is modeled as $p_{\theta}(\bm{x_{0:T}}):=p(\bm{x_{T}})\prod_{t=1}^{T}p_{\theta}(\bm{x_{t-1}}\,|\,\bm{x_{t}})$, and is trained by maximizing the evidence lower bound (ELBO) [15, 5]: $\mathbb{E}q\left[\log\frac{p_{\theta}(\bm{x_{0:T}})}{q(\bm{x_{1:T}}\mid\bm{x_{0}})}\right]$. After training, samples can be generated by first drawing $\bm{x_{T}}\sim p(\bm{x_{T}})$ and then sequentially applying the learned reverse transitions to obtain $\bm{x_{0}}$. For conditional generation tasks, the reverse process can be extended to model $p_{\theta}(\bm{x_{t-1}}\,|\,\bm{x_{t}},c)$, where $c$ denotes the conditioning information.
Expectile Regression. Expectile regression has been extensively studied in econometrics [26] and recently introduced in offline RL [19]. The $\tau$-expectile (with $\tau\in(0,1)$) of a real-valued random variable $\bm{X}$ is defined as the solution to the asymmetric least squares problem:
@@ 4 Advantage-Based Diffusion Actor-Critic Algorithm
Diffusion Policy. We model our policy as the reverse process of a conditional diffusion model [36]:
@@
where the terminal sample $a^{0}$ is used as the action for RL evaluation. During training, we sample $(\bm{s},\bm{a})$ pairs from the offline dataset $\mathcal{D}$ and construct noisy samples $\bm{a^{i}}=\sqrt{\bar{\alpha}_{i}}\bm{a}+\sqrt{1-\bar{\alpha}_{i}}\bm{\epsilon}$ (Eq. (3)), where $\alpha_{i}=1-\beta_{i}$, $\bar{\alpha}_{i}=\prod_{j=1}^{i}\alpha_{j}$, and $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. Following DDPM [13], we train the following noise prediction model $\bm{\epsilon}_{\theta}(\bm{a^{i}},\bm{s},i)$ to approximate the added noise, which determines the reverse process $p_{\theta}(\bm{a^{i-1}}\,|\,\bm{a^{i}},\bm{s})$:
@@
We summarize our implementation in Algorithm 1. A key innovation of our method lies in the use of $A(\bm{a}\,|\,\bm{s})$; without this component, our approach degenerates to the DQL method [36], which uses standard Bellman backup. In addition to achieving new SOTA, we compare with DQL on PointMaze in Section 5.2 to visualize how advantage modulation facilitates superior OOD action selection.
@@ 5 Experiments
In this section, we begin by evaluating our method on the widely recognized D4RL benchmark [9]. We then design a dedicated experiment on PointMaze—another task within D4RL—to empirically and visually demonstrate ADAC’s ability to identify beneficial OOD actions. Finally, we perform an ablation study to dissect the contribution of key components in our method.
@@
Baseline. We consider a diverse array of baseline methods that exhibits strong results in each domain of tasks. For policy regularization-based method, we compare with the classic BC, TD3+BC [10], BEAR[20], BRAC[37], BCQ[11], AWR[27], O-RL[6], and DQL[36]. For pessimistic value function-based approach, we include CQL[21], IQL[19], and REM[1]. For model-based offline RL, we choose MoRel[18]. For the classic online method, we include SAC[12]. For conditional sequence modeling approaches, we include DT[7], Diffuser[14], and DD[2]. We report the performance of baseline methods either from the best results published in their respective papers or from [36].
@@ 5.1 Benchmark Results
Results for Gym Tasks. While prior methods already achieve strong performance on Gym tasks, ADAC consistently delivers further improvements. Among all Gym tasks, halfcheetah is notably the most challenging, often yielding the lowest scores. However, ADAC achieves the largest performance boost on this task family, with an improvement of 10%. According to the protocol established in [9], a normalized score of 100 corresponds to an expert-level policy. Therefore, achieving an average normalized score exceeding 90 demonstrates the practical effectiveness of our method.
@@ 5.2 Visualizing OOD Action Selection in PointMaze
We train both DQL [36] and ADAC for 50 000 steps using the constructed dataset to obtain their respective policies. Each policy is then evaluated in our environment, generating 300 trajectories per method, as illustrated in the bottom row of Figure 4.
@@ 5.3 Ablation Study
One potential limitation of our method is the need to sample multiple actions ($N\equiv 25$ in this paper) from the behavior policy to compute the advantage function. However, our ablation study shows that in each task domain, competitive performance can be achieved across at least three different values of $\kappa$, indicating that the algorithm is relatively insensitive to $\kappa$. This suggests that the number of candidate actions can be reduced without significantly affecting performance. We implement our algorithm using the JAX/Flax framework, which offers faster training and inference speed than the DQL method and is comparable to the optimized EDP [17] implementation as shown in Figure 9.

### v3

@@ Abstract
Offline reinforcement learning (RL) aims to learn decision-making policies from fixed datasets without online interactions, providing a practical solution where online data collection is expensive or risky. However, offline RL often suffers from distribution shift, resulting in inaccurate evaluation and substantial overestimation on out-of-distribution (OOD) actions. To address this, existing approaches incorporate conservatism by indiscriminately discouraging all OOD actions, thereby hindering the agent’s ability to generalize and exploit beneficial ones. In this paper, we propose Advantage-based Diffusion Actor-Critic (ADAC), a novel method that systematically evaluates OOD actions using the batch-optimal value function. Based on this evaluation, ADAC defines an advantage function to modulate the Q-function update, enabling more precise assessment of OOD action quality. We design a custom PointMaze environment and collect datasets to visually reveal that advantage modulation can effectively identify and select superior OOD actions. Extensive experiments show that ADAC achieves state-of-the-art performance on almost all tasks in the D4RL benchmark, with particularly clear margins on the more challenging tasks. Our code is available at https://github.com/NUS-CORE/adac.
@@ 1 Introduction
Offline reinforcement learning (RL) lange2012batch ; levine2020offline focuses on learning decision-making policies solely from previously collected datasets, entirely circumventing the need for online interactions with the environment. This paradigm is particularly appealing for applications where online data collection is prohibitively expensive or poses safety concerns, thereby fueling its growing adoption kalashnikov2018scalable ; prudencio2023survey . While conceptually attractive, offline RL often suffers from distribution shift between the behavior policy and the learned policy. The policy evaluation on out-of-distribution (OOD) actions is prone to extrapolation error fujimoto2019off , which can be amplified through bootstrapping, leading to significant overestimation kumar2020conservative .
To mitigate overestimation, a common strategy in offline RL is to incorporate conservatism into algorithm design. Value-based methods kumar2020conservative ; kostrikov2021offline ; lyu2022mildly achieve this by learning a pessimistic value function that reduces the estimated value of OOD actions to discourage their selection. Alternatively, policy-based methods fujimoto2021minimalist ; fujimoto2019off ; wang2022diffusion enforce conservatism by constraining the learned policy to remain close to the behavior policy, thereby avoiding querying OOD actions. Similarly, conditional sequence modeling approaches chen2021decision ; janner2022planning ; ajay2022conditional inherently induce conservative behavior by restricting the policy to imitate the behavior contained in the offline dataset. In a distinct manner, model-based approaches kidambi2020morel ; yu2020mopo ; sun2023model ensure conservatism by learning a pessimistic dynamics model where uncertainty-based penalization systematically underestimates the value of OOD actions.
@@
To this end, we propose Advantage-based Diffusion Actor-Critic (ADAC), a novel method that systematically assesses the quality of OOD actions, selectively encourages beneficial ones, and discourages risky ones to maintain conservatism. Our key idea builds on the insight that the optimal value function provides a natural baseline for evaluating state desirability. Specifically, we regard an OOD action as advantageous if it can move the current state to a successor state whose value, under the optimal value function, exceeds that of any reachable state under the behavior policy. Since the true optimal value function is inaccessible from offline data, we adopt the batch-optimal value function ($V_{\mu}^{\ast}$ in Figure 1) as a principled surrogate. We show theoretically that the batch-optimal value function can be reliably approximated via expectile regression on the dataset. Based on this approximation, we define an advantage function to assess the desirability of actions, which is then used to modulate the temporal difference (TD) target during Q-function (critic) learning. The policy (actor) is parameterized as a diffusion model due to its strong expressiveness ho2020denoising ; wang2022diffusion and is updated under the guidance of the learned Q-function. We summarize the full architecture of ADAC in Figure 1. To the best of our knowledge, ADAC represents a pioneering attempt to systematically evaluate OOD actions, enabling the selection of beneficial ones. As a result, ADAC achieves state-of-the-art (SOTA) performance on the majority of the D4RL fu2020d4rl benchmark tasks.
@@ 2 Preliminaries
Offline Reinforcement Learning. RL problems are commonly formulated within the framework of a Markov Decision Process (MDP), defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,\rho_{0},P,\gamma)$. Here, $\mathcal{S}$ denotes the state space, $\mathcal{A}$ represents the action space, $r(\bm{s},\bm{a}):\mathcal{S}\times\mathcal{A}\to[-R_{\max},R_{\max}]$ is a bounded reward function, $\rho_{0}(\bm{s})$ specifies the initial state distribution, $P(\bm{s^{\prime}}\,|\,\bm{s},\bm{a}):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}_{+}$ defines the transition dynamics, and $\gamma\in(0,1)$ is the discount factor sutton2018reinforcement .
@@
Diffusion Model. Diffusion models sohl2015deep ; ho2020denoising ; song2020score consist of a forward process that corrupts data with noise and a reverse process that reconstructs data from noise. Specifically, the forward process is conducted by gradually adding Gaussian noise to samples $\bm{x_{0}}$ from an unknown data distribution $p_{\theta}(\bm{x_{0}})$, formulated as:
@@
where $T$ denotes the total number of diffusion steps, and $\alpha_{t}$ controls the variance of the added noise at each step $t$. The reverse process is modeled as $p_{\theta}(\bm{x_{0:T}}):=p(\bm{x_{T}})\prod_{t=1}^{T}p_{\theta}(\bm{x_{t-1}}\,|\,\bm{x_{t}})$, and is trained by maximizing the evidence lower bound (ELBO) jordan1999introduction ; blei2017variational : $\mathbb{E}q\left[\log\frac{p_{\theta}(\bm{x_{0:T}})}{q(\bm{x_{1:T}}\mid\bm{x_{0}})}\right]$. After training, samples can be generated by first drawing $\bm{x_{T}}\sim p(\bm{x_{T}})$ and then sequentially applying the learned reverse transitions to obtain $\bm{x_{0}}$. For conditional generation tasks, the reverse process can be extended to model $p_{\theta}(\bm{x_{t-1}}\,|\,\bm{x_{t}},c)$, where $c$ denotes the conditioning information.
Expectile Regression. Expectile regression has been extensively studied in econometrics newey1987asymmetric and recently introduced in offline RL kostrikov2021offline . The $\tau$-expectile (with $\tau\in(0,1)$) of a real-valued random variable $\bm{X}$ is defined as the solution to the asymmetric least squares problem:
@@ 4 Advantage-Based Diffusion Actor-Critic Algorithm
Diffusion Policy. We model our policy as the reverse process of a conditional diffusion model wang2022diffusion :
@@
where the terminal sample $a^{0}$ is used as the action for RL evaluation. During training, we sample $(\bm{s},\bm{a})$ pairs from the offline dataset $\mathcal{D}$ and construct noisy samples $\bm{a^{i}}=\sqrt{\bar{\alpha}_{i}}\bm{a}+\sqrt{1-\bar{\alpha}_{i}}\bm{\epsilon}$ (Eq. (3)), where $\alpha_{i}=1-\beta_{i}$, $\bar{\alpha}_{i}=\prod_{j=1}^{i}\alpha_{j}$, and $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. Following DDPM ho2020denoising , we train the following noise prediction model $\bm{\epsilon}_{\theta}(\bm{a^{i}},\bm{s},i)$ to approximate the added noise, which determines the reverse process $p_{\theta}(\bm{a^{i-1}}\,|\,\bm{a^{i}},\bm{s})$:
@@
We summarize our implementation in Algorithm 1. A key innovation of our method lies in the use of $A(\bm{a}\,|\,\bm{s})$; without this component, our approach degenerates to the DQL method wang2022diffusion , which uses standard Bellman backup. In addition to achieving new SOTA, we compare with DQL on PointMaze in Section 5.2 to visualize how advantage modulation facilitates superior OOD action selection.
@@ 5 Experiments
In this section, we begin by evaluating our method on the widely recognized D4RL benchmark fu2020d4rl . We then design a dedicated experiment on PointMaze—another task within D4RL—to empirically and visually demonstrate ADAC’s ability to identify beneficial OOD actions. Finally, we perform an ablation study to dissect the contribution of key components in our method.
@@
Baseline. We consider a diverse array of baseline methods that exhibits strong results in each domain of tasks. For policy regularization-based method, we compare with the classic BC, TD3+BC fujimoto2021minimalist , BEARkumar2019stabilizing , BRACwu2019behavior , BCQfujimoto2019off , AWRpeng2019advantage , O-RLbrandfonbrener2021offline , and DQLwang2022diffusion . For pessimistic value function-based approach, we include CQLkumar2020conservative , IQLkostrikov2021offline , and REMagarwal2020optimistic . For model-based offline RL, we choose MoRelkidambi2020morel . For the classic online method, we include SAChaarnoja2018soft . For conditional sequence modeling approaches, we include DTchen2021decision , Diffuserjanner2022planning , and DDajay2022conditional . We report the performance of baseline methods either from the best results published in their respective papers or from wang2022diffusion .
@@ 5.1 Benchmark Results
Results for Gym Tasks. While prior methods already achieve strong performance on Gym tasks, ADAC consistently delivers further improvements. Among all Gym tasks, halfcheetah is notably the most challenging, often yielding the lowest scores. However, ADAC achieves the largest performance boost on this task family, with an improvement of 10%. According to the protocol established in fu2020d4rl , a normalized score of 100 corresponds to an expert-level policy. Therefore, achieving an average normalized score exceeding 90 demonstrates the practical effectiveness of our method.
@@ 5.2 Visualizing OOD Action Selection in PointMaze
We train both DQL wang2022diffusion and ADAC for 50 000 steps using the constructed dataset to obtain their respective policies. Each policy is then evaluated in our environment, generating 300 trajectories per method, as illustrated in the bottom row of Figure 4.
@@ 5.3 Ablation Study
One potential limitation of our method is the need to sample multiple actions ($N\equiv 25$ in this paper) from the behavior policy to compute the advantage function. However, our ablation study shows that in each task domain, competitive performance can be achieved across at least three different values of $\kappa$, indicating that the algorithm is relatively insensitive to $\kappa$. This suggests that the number of candidate actions can be reduced without significantly affecting performance. We implement our algorithm using the JAX/Flax framework, which offers faster training and inference speed than the DQL method and is comparable to the optimized EDP kang2023efficient implementation as shown in Figure 9.